# A Mathematical Theory for Learning Semantic Languages by Abstract Learners

### v2

@@ V Hierarchy of skills
In view of (24), there are two thresholds for achieving low testing errors in a domain-specific task: (i) a large number of basic skills are learned when the number of basic texts exceeds the threshold in the foundation model, and (ii) a large number of domain-specific skills are learned when the number of domain-specific texts exceeds the threshold in the fine-tuning model. In Figure 9, we present a contour plot of the probability $p_{t}$ as a function of $R$ and $R_{f}$ for the 1-skill learner with values of $c=3$ and $c_{f}=3$, and the generating function $\Lambda_{f}(x)=e^{-3(1-x)}$, i.e., the number of basic skills required by a domain-specific skill is Poisson distributed with mean $3$. The graph features two percolation thresholds associated with $R$ and $R_{f}$. It is observed that if $R<1$, the majority of domain-specific text nodes remain unlearned, primarily because the basic skills are not acquired, indicating the first percolation threshold. Conversely, when $R>1$, attention shifts to the second percolation threshold, $R_{f}$; in this scenario, a significant number of domain-specific texts are learned, provided that $R_{f}>1$. This distinction underscores the critical values of $R$ and $R_{f}$ in determining the learning outcomes within the graph.
VI Application to Semantic Compression
Traditionally, lossless compression of a text requires exact recovery of the sequence of tokens in a text. However, if our interest lies in recovering only the semantic meaning of a text, we might be able to compress it using fewer bits than required for lossless compression. This concept is referred to as the semantic (level B) problem in [29].
Definition 7
Here, we demonstrate how to use generative learners for semantic compression. Once the training of an abstract learner is complete, the expected number of skills learned is $|S|(1-e^{-cR(1-p)})$ with $p$ given by (3) for the 1-skill learner and by (7) for the Poisson learner. By indexing the learned skills, the number of bits required to represent a learned skill is $\log_{2}(|S|(1-e^{-cR(1-p)}))$. Therefore, if a text with $k$ skills is understood by the learner, it can be compressed/encoded using $k\log_{2}(|S|(1-e^{-cR(1-p)}))$ bits. As previously shown, the probability that a randomly selected text is understood by the learner is $e^{-(ce^{-cR(1-p)})}$, and it requires on average $c\log_{2}(|S|(1-e^{-cR(1-p)}))$ bits to encode the text. Conversely, if a randomly selected text is not understood by the learner, it can be encoded using a lossless compression encoder. Suppose the lossless compression encoder requires, on average, $z$ bits to compress a text. Then the semantic compression method described above requires, on average,
$e^{-(ce^{-cR(1-p)})}c\log_{2}(|S|(1-e^{-cR(1-p)}))+(1-e^{-(ce^{-cR(1-p)})})z$
bits for a text. By Shannon’s analysis in [30], the entropy per word in the English language is approximately $11.82$ bits. By assuming that an average sentence has approximately $20$ words, the number of bits required to represent a text is approximately $z=236.4$ bits. With $R$ sufficiently large such that $c\log_{2}(|S|(1-e^{-cR(1-p)}))\approx c\log_{2}|S|$ and with $c$ set as $5$, a compression gain is obtained as long as $|S|\leq 2^{47.28}$.
For semantic communication, we can utilize a semantic encoder/decoder. In Figure 10, we present the general architecture of a semantic communication system, which consists of a semantic transmitter, a semantic receiver, and a physical channel. The semantic transmitter includes a semantic encoder and a channel encoder, while the semantic receiver includes a semantic decoder and a channel decoder. The communication between the channel encoder and decoder occurs over the physical channel, referred to as level A communication in [29]. Conversely, the communication between the semantic encoder and decoder is not over the physical channel and is termed level B communication in [29].
Recent literature on semantic communications, such as [31, 32, 33], has proposed an end-to-end approach for jointly training the semantic and channel encoder/decoder. This method is claimed to be superior to separate training. However, the end-to-end approach does not scale efficiently with data size. As demonstrated in the previous section, a large dataset is necessary for training to exhibit the emergence of semantic capability. In light of this, a modular design akin to Internet protocol may be more effective for semantic communication.
Figure 10: The general framework of a semantic communication system.
For domain-specific texts discussed in Section V, semantic compression can also be used once fine-tuning is complete. If the number of domain-specific skills is much smaller than the number of basic skills, i.e., $|S_{f}|\ll|S|$, one expects the compression gain for domain-specific texts in the fine-tuning model will surpass that achieved using the foundation model.
VII Extensions to multiple classes of skills and texts
In this section, we extend the analysis to the settings with multiple classes of skills and texts. The motivation for such an extension is the existence of multiple subjects in texts, such as math, physics, chemistry, law, etc.
VII-A Poisson learners with multiple classes of skills and texts
@@
VII-A1 The ensemble of random bipartite graphs
Suppose that the set of texts $T$ can be classified into $J$ disjoint classes of texts, $T_{1},T_{2},\ldots,T_{J}$, and that the set of skills $S$ can be classified into $K$ disjoint classes of skills, $S_{1},S_{2},\ldots,S_{K}$. As before, we sample a subset of $D$ texts from the set of texts $T$ and present these $D$ texts as the training texts. Suppose that there are $D_{j}$ class $j$ texts, $j=1,2,\ldots,J$. Let $R=D/|S|$ be the ratio of the number of training texts to the number of skills. Also, let $\alpha_{j}=D_{j}/D$, $j=1,2,\ldots,J$ be the ratio of the number of class $j$ training texts to the total number of training texts, and $\beta_{k}=|S_{k}|/|S|$, $k=1,2,\ldots,K$, be the ratio of the number of class $k$ skills to the total number of skills. As before, the skill-text bipartite graph of a sampled semantic language is a random bipartite graph with $|S|$ skill nodes on one side and $D$ text nodes on the other side. An edge exists between a skill node $s$ and a text node $t$ if $s\in\phi(t)$.
We make the following two assumptions about a randomly selected text:
(Poisson distribution) For a randomly selected class $j$ text $t\in T_{j}$, the number of class $k$ skills in $t$, for $k=1,2,\ldots,K$, are independent Poisson random variables with means $c_{k,j}$, for each $k=1,2,\ldots,K$ and $j=1,2,\ldots,J$.
(Uniform distribution) An edge is called a class $(k,j)$-edge if it connects a class $k$ skill node to a class $j$ text node. The probability that a class $(k,j)$-edge of the (randomly selected) text node $t\in T_{j}$ is connected to a particular skill node $s\in S_{k}$ is $1/|S_{k}|$.
From (A1m) and (A2m), the number of edges between a class $j$ text node and a class $k$ skill node follows a Poisson distribution with mean $c_{k,j}/|S_{k}|$. Since there are $\alpha_{j}D$ class $j$ text nodes, the number of edges between class $k$ skill nodes and class $j$ text nodes is also Poisson distributed with mean $c_{k,j}\alpha_{j}D/|S_{k}|=c_{k,j}\frac{\alpha_{j}}{\beta_{k}}R$. This leads to the following proposition.
@@ VII-A1 The ensemble of random bipartite graphs
Consider the sampled skill-text bipartite graph from a semantic language ${\cal L}=({\cal A},T,S,\phi)$ with multiple classes of skills and texts. Then:
(i) (Poisson distribution) For a randomly selected class $k$ skill node $s$, the number of class $j$ text nodes connected to $s$, for $j=1,2,\ldots,J$, are independent Poisson random variables with means $d_{k,j}$, for each $k=1,2,\ldots,K$ and $j=1,2,\ldots,J$, where
$d_{k,j}=c_{k,j}\frac{\alpha_{j}}{\beta_{k}}R.$ (25)
As such, the degree of a class $k$ skill node follows a Poisson distribution with mean $d_{k}$, where
$d_{k}=\sum_{j=1}^{J}c_{k,j}\frac{\alpha_{j}}{\beta_{k}}R.$ (26)
(ii) (Uniform distribution) The probability that a class $(k,j)$-edge of a randomly selected skill node $s$ is connected to a particular class $j$ text node is $1/(\alpha_{j}D)$.
@@
Call an edge a class $k$ edge if the skill end of the edge is connected to a class $k$ skill. From Proposition 8, the probability that a class $k$ edge is a class $(k,j)$-edge is
@@
VII-A2 Poisson learners with multiple classes of skills
We say a Poisson learner with $K$ classes of skills is subject to a Poisson offered load $\rho=(\rho_{1},\rho_{2},\ldots,\rho_{K})$ from a text $t$ if the number of class $k$ skills connected to the text node $t$, $k=1,2,\ldots,K$, are independent Poisson random variables with means $\rho_{k}$, $k=1,2,\ldots,K$. Analogous to the definition of a Poisson receiver with multiple class of input traffic in [27], we define a Poisson learner with multiple class of skills below.
@@ VII-A2 Poisson learners with multiple classes of skills
VII-A3 Density evolution analysis
Now we repeatedly present the $J$ classes of texts (a total of $D$ texts) to a Poisson learner with $K$ classes of skills to learn the skills. Our study utilizes the tree evaluation approach, as detailed in [23, 17, 34, 19], along with the concepts of reduced Poisson offered load, referenced in [27]. The tree evaluation technique, further developed in [24, 25], has been instrumental in determining the capacity of low-density parity-check (LDPC) codes [14]. This method is popularly recognized as the density evolution analysis within the field of information theory. Additionally, the reduced Poisson offered load approach is a prominent method for analyzing queueing networks, as exemplified by references such as [35, 36, 37].
Similar to the analysis in the original IRSA [17] and the coded Poisson receivers [26], we define $\Lambda_{k,\ell}$ as the probability that a class $k$ skill node has $\ell$ edges. From Proposition 8, we know that
@@ VII-A3 Density evolution analysis
Our density evolution analysis consists of the following steps:
@@
We can represent the offered load at a class $j$ Poisson receiver after the $i^{th}$ SCNS iteration by the vector $q^{(i)}\circ\tilde{c}_{j}$, where $\circ$ denotes the element-wise multiplication of two vectors.
@@
That (35) holds follows directly from the definition of a Poisson learner in Definition 9 as the offered load of a class $j$ text to the Poisson learner after the $i^{th}$ SCNS iteration is $q^{(i)}\circ\tilde{c}_{j}$.
@@
$q_{k}^{(i+1)}=\lambda_{k}(1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i)}\circ\tilde{c}_{j})),$ (37)
@@
(v) Let $\tilde{P}_{{\rm suc},k}^{(i)}$ be the probability that a randomly selected class $k$ skill can be learned after the $i^{th}$ iteration. Such a probability is the probability that at least one text end of the edges connected to $s$ has been learned after the $i^{th}$ iteration. Since the probability that a randomly selected class $k$ skill node has $\ell$ edges is $\Lambda_{k,\ell}$, we have from (36) that
@@
We note that here we only outline the steps to derive the recursive equations. A rigorous proof for the density evolution analysis requires the concentration theorem (Theorem 2 of [25]) to prove the ergodicity of the system. i.e., as $D\to\infty$, the average fraction of skill nodes that have not been learned after the $i^{th}$ SCNS iteration converges to the probability that a randomly selected skill node has not been learned after the $i^{th}$ SCNS iteration.
@@
$k=1,2,\ldots,K$. These $K$ coupled equations can be used for obtaining the scaling law for $p$ with respect to $R$, given the set of parameters $c_{k,j}$’s, $\alpha_{j}$’s, and $\beta_{k}$’s. Note that (42) reduces to (7) when $J=K=1$ and $i\to\infty$.
VII-B Deterministic $\psi$-learners
@@
One typical example of the $\psi$-learner is 1-skill learner. As before, one can also generalize 1-skill learner to $D$-skill learner, where there are at most $D$ novel skills can be learned. The $D$-skill learner is a $\psi$-learner with
@@
Analogous to the proof of Theorem 14 in [28], it is shown that for every $\psi$-learner, there is an induced Poisson learner. This relationship is established by computing the throughput of the $\psi$-learner when it is subject to a Poisson offered load $\rho$, and then using (28) to determine the success probability function. For the $D$-skill learner, the throughput for a Poisson offered load $\rho$ is given by
@@
Figure 12: The plot illustrates the computed values of $p_{1}$ (and $p_{2}$) for the near-far learner as a function of $R$, which ranges from 0.1 to 10 with a step size of 0.0.5, for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. The computation is based on the two recursive equations in (VII-B) with the initial condition $p_{1}^{(0)}=p_{2}^{(0)}=1$, where $P_{{\rm suc},k}(\rho)$, $k=1$ and 2, are defined in (VII-B). The number of iterations is 100.
@@
In Figure 12, we plot the computed values of $p_{1}$ as a function of $R$ for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. For this example, it is noted that the values of $p_{2}$ overlap with those of $p_{1}$, making $p_{2}$ indistinguishable in the figure. Interestingly, thresholds are also observed for these five curves. As demonstrated in [26], the thresholds for these two classes of skills coincide due to the coupling of the density evolution equations (similar to those in (VII-B)), leading to the simultaneous emergence of various classes of skills.
VIII Discussions and Future Directions
@@
Inspired by [13], we defined a semantic language as a skill-text bipartite graph. Skills can be learned by repeatedly presenting training texts to abstract learners, such as 1-skill learners and Poisson learners. By using density evolution analysis, we showed the emergence of learned skills when the ratio $R$ exceeds a certain threshold. Moreover, we derived the corresponding scaling law for testing errors. We also demonstrated how trained abstract learners can be used for semantic compression and semantic communication.
@@
The present work is intended to inspire further thoughts on the emergence of learned skills in today’s LLMs, drawing relation to the iterative decoding procedure in communications and information theory. However, we recognize that there are still significant gaps between our theory and real-world systems. Several key factors that should be further investigated to better link theory with practice are listed below:
@@
    Scaling laws of the number of parameters: The scaling law obtained in this paper only pertains to the size of training texts. It is unclear how testing errors scale with the number of parameters. For instance, GPT-3 is known to have more than 1,750 billion parameters, and LLMs prior to GPT-3 do not exhibit the emergence of skills. One possible explanation for the necessity of a larger number of parameters is that the capacity of a transformer, defined as the number of skills (or patterns) that can be stored in the corresponding Hopfield network [12], depends on the number of parameters in the transformer.
@@
$D$ The number of training texts
$D_{f}$ The number of domain-specific texts for training
$D_{j}$ The number of class $j$ texts
$d_{k,j}$ The average number of class $j$ texts connected to a class $k$ skill node (see (25))
$d_{k}$ The mean degree of a class $k$ skill node (see (26))
@@
$p_{s}$ The probability that a randomly selected skill is learned
@@
$p_{t}$ The probability of testing errors of a randomly selected text from the domain-specific texts
@@
$S$ The set of skills
$S_{k}$ The set of class $k$ skills
$S_{f}$ The set of domain-specific skills
$T$ The set of texts composed of sequences of tokens
$T_{j}$ The set of class $j$ texts
@@
$\mu_{s}$ The probability that a skill node is not connected to the giant component via one of its edges
$\mu_{t}$ The probability that a text node is not connected to the giant component via one of its edges

### v3

@@ V Hierarchy of skills
In view of (26), there are two thresholds for achieving low testing errors in a domain-specific task: (i) a large number of basic skills are learned when the number of basic texts exceeds the threshold in the foundation model, and (ii) a large number of domain-specific skills are learned when the number of domain-specific texts exceeds the threshold in the fine-tuning model. In Figure 11, we present a contour plot of the testing error $\epsilon_{f}$ as a function of $R$ and $R_{f}$ for the 1-skill learner with values of $c=3$ and $c_{f}=3$. The number of basic skills required by a domain-specific skill is Poisson distributed with mean $3$, in which case $\Lambda_{f}(x)=e^{-3(1-x)}$. The graph features two percolation thresholds associated with $R$ and $R_{f}$. It is observed that if $R<1$, the majority of domain-specific text nodes remain unlearned, primarily because the basic skills are not acquired, indicating the first percolation threshold. Conversely, when $R>1$, attention shifts to the second percolation threshold, $R_{f}$; in this scenario, a significant number of domain-specific texts are learned, provided that $R_{f}>1$. This distinction underscores the critical values of $R$ and $R_{f}$ in determining the learning outcomes within the graph.
VI Extensions to multiple classes of skills and texts
In this section, we extend the analysis to the setting with multiple classes of skills and texts, which is motivated by the existence of multiple subjects in texts, such as math, physics, chemistry, law, etc. Note that, in these settings, a text may require skills from multiple classes to understand and, thus, introduces interdependence among classes. However, different from Section V, the multiple classes of skills are not assumed to be hierarchical in nature and, thus, could be learned without a prerequisite on the knowledge of other classes of skills.
VI-A Poisson learners with multiple classes of skills/texts
@@
VI-A1 The ensemble of random bipartite graphs
[…]
We make the following two assumptions about the sampled texts:
(Poisson degree distribution) Denote by $N_{k,j}(t)$ the number of class $k$ skills in a text $t$ in the set of class $j$ texts ${\cal D}_{j}$. Then $\{N_{k,j}(t),t\in{\cal D}_{j}\}$, $j=1,2,\ldots,J$, $k=1,2,\ldots,K$, are independent Poisson random variables with means $c_{k,j}$.
(Uniform connection) Each class $(k,j)$-edge of a text node $t\in{\cal D}_{j}$ is connected to a skill node in ${\cal S}_{k}$ independently and uniformly. Thus, the probability that a class $(k,j)$-edge from a text $t\in{\cal D}_{j}$ is connected to a particular skill node $s\in{\cal S}_{k}$ is $1/|{\cal S}_{k}|$.
[…]
@@ VI-A1 The ensemble of random bipartite graphs
[…]
(ii) (Uniform connection) Each class $(k,j)$-edge of a class $k$ skill node is connected to a class $j$ text node independently and uniformly. Thus, the probability that a class $(k,j)$-edge from a skill node $s\in{\cal S}_{k}$ is connected to a particular class $j$ text node is $1/(\alpha_{j}|{\cal D}|)$.
@@
An edge is a class $k$ edge if the skill end of the edge is connected to a class $k$ skill. From Proposition 7, the probability that a class $k$ edge is a class $(k,j)$-edge is
@@
VI-A2 Density evolution for Poisson learners with multiple classes of skills
We say a Poisson learner with $K$ classes of skills is subject to a Poisson offered load $\rho=(\rho_{1},\rho_{2},\ldots,\rho_{K})$ from a text $t$ if the number of class $k$ skills connected to the text node $t$, $k=1,2,\ldots,K$, are independent Poisson random variables with means $\rho_{k}$, $k=1,2,\ldots,K$. Analogous to the definition of a Poisson receiver with multiple class of input traffic in [27], we can define a Poisson learner with multiple class of skills as follows.
@@ VI-A2 Density evolution for Poisson learners with multiple classes of skills
Suppose that $J$ classes of texts (a total of $|{\cal D}|$ texts) are presented to a Poisson learner with $K$ classes of skills. Our study utilizes the tree evaluation approach, as detailed in [23, 17, 29, 19], along with the concepts of reduced Poisson offered load, referenced in [27]. The tree evaluation method is also recognized as the density evolution analysis within the field of information theory [14, 24, 25]. Similar to the analysis in the original IRSA [17] and the coded Poisson receivers [26], we define $\Lambda_{k,\ell}$ as the probability that a class $k$ skill node has $\ell$ edges. From Proposition 7, we know that
@@
Our density evolution analysis involves the following steps:
@@
We can represent the offered load at a class $j$ Poisson receiver after the $i^{th}$ SCNS iteration by the vector $q^{(i)}\circ\tilde{c}_{j}$, where $\circ$ denotes the element-wise multiplication of two vectors. This reduced load argument is also the key step that greatly reduces the computational complexity in the density evolution method.
@@
The above equation follows directly from the definition of a Poisson learner in Definition 8 as the offered load of a class $j$ text to the Poisson learner after the $i^{th}$ SCNS iteration is $q^{(i)}\circ\tilde{c}_{j}$.
@@
$q_{k}^{(i+1)}=\lambda_{k}\Big{(}1-\sum_{j=1}^{J}r_{k,j}P_{{\rm suc},k}(q^{(i)}\circ\tilde{c}_{j})\Big{)},$ (37)
@@
(v) Let $\tilde{P}_{{\rm suc},k}^{(i)}$ be the probability that a randomly selected class $k$ skill can be learned after the $i^{th}$ iteration. This is the probability that at least one text end of the edges connected to $s$ has been learned after the $i^{th}$ iteration. Since the probability that a randomly selected class $k$ skill node has $\ell$ edges is $\Lambda_{k,\ell}$, we have from (36) that
@@
Here, we only outline the steps to derive the recursive equations. A rigorous proof for the density evolution analysis requires the concentration theorem (Theorem 2 of [25]) to prove the ergodicity of the system. i.e., as $|{\cal D}|\to\infty$, the average fraction of skill nodes that have not been learned after the $i^{th}$ SCNS iteration converges to the probability that a randomly selected skill node has not been learned after the $i^{th}$ SCNS iteration.
@@
$k=1,2,\ldots,K$. These $K$ coupled equations can be used for obtaining the scaling law for $p$ with respect to $R$, given the parameters $c_{k,j}$’s, $\alpha_{j}$’s, and $\beta_{k}$’s. Note that (41) reduces to (8) when $J=K=1$ and $i\to\infty$.
VI-B Deterministic $\psi$-learners
@@
One typical example of the $\psi$-learner is 1-skill learner. As before, one can also generalize 1-skill learner to $|{\cal D}|$-skill learner, where there are at most $|{\cal D}|$ novel skills can be learned. The $|{\cal D}|$-skill learner is a $\psi$-learner with
@@
Analogous to the proof of Theorem 14 in [28], it is shown that for every $\psi$-learner, there is an induced Poisson learner. This relationship is established by computing the throughput of the $\psi$-learner when it is subject to a Poisson offered load $\rho$, and then using (28) to determine the success probability function. For the $|{\cal D}|$-skill learner, the throughput for a Poisson offered load $\rho$ is given by
@@
Figure 13: The plot illustrates the computed values of $p_{1}$ (and $p_{2}$) for the near-far learner as a function of $R$, which ranges from 0.1 to 10 with a step size of 0.05, for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. The computation is based on the two recursive equations in (VI-B) with the initial condition $p_{1}^{(0)}=p_{2}^{(0)}=1$, where $P_{{\rm suc},k}(\rho)$, $k=1$ and 2, are defined in (VI-B). The number of iterations is 100.
@@
In Figure 13, we plot the computed values of $p_{1}$ as a function of $R$ for each $\beta$ value (0.1, 0.3, 0.5, 0.7, 0.9). The average degrees of the two classes of skills are $c_{1}=3$ and $c_{2}=7$, respectively. For this example, it is noted that the values of $p_{2}$ overlap with those of $p_{1}$, making $p_{2}$ indistinguishable in the figure. Interestingly, thresholds are also observed for these five curves. As demonstrated in [26], the thresholds for these two classes of skills coincide due to the coupling of the density evolution equations (similar to those in (VI-B)), leading to the simultaneous emergence of various classes of skills.
VII Application to Semantic Compression and Communication
The lossless compression of a text requires exact recovery of the sequence of tokens in a text. However, if our interest lies in recovering only the semantic meaning of a text, we might be able to compress it using fewer bits than required for lossless compression. This concept is referred to as the semantic (level B) problem in [31].
Definition 10
[…]
For semantic communication, we can utilize a semantic encoder/decoder. In Figure 14, we present the general architecture of a semantic communication system, which consists of a semantic transmitter, a semantic receiver, and a physical channel. The semantic transmitter includes a semantic encoder and a channel encoder, while the semantic receiver includes a semantic decoder and a channel decoder. The communication between the channel encoder and decoder occurs over the physical channel, referred to as level A communication in [31]. Conversely, the communication between the semantic encoder and decoder is termed level B communication in [31].
Recent literature on semantic communications, such as [33, 34, 35], has proposed an end-to-end approach for jointly training the semantic and channel encoder/decoder. This method is claimed to be superior to separate training. However, the end-to-end approach does not scale efficiently with data size. In fact, as shown in the previous section, a large dataset is necessary for training to exhibit the emergence of semantic capability. Hence, for the transmission of texts from a general semantic language, an LLM model is required at both the semantic transmitter and receiver, which would be difficult to retrain and adapt to varying physical channels. In light of this, a modular design (i.e., separate source and channel coding) may be more effective for semantic communication in practice.
Figure 14: The general framework of a semantic communication system.
For domain-specific texts discussed in Section V, semantic compression can also be used once fine-tuning is complete. If the number of domain-specific skills is much smaller than the number of basic skills, i.e., |$𝒮|\color[rgb]{0,0,0}{\cal S}_{f}|\ll|{\cal S}|$, one expects the compression gain for domain-specific texts in the fine-tuning model will surpass that achieved using the foundation model.
VIII Conclusions and Future Directions
@@
Inspired by [13], we defined a semantic language as a skill-text bipartite graph. Skills can be learned by repeatedly presenting training texts to abstract learners, such as 1-skill learners and Poisson learners. By using density evolution analysis, we showed the emergence of learned skills when the ratio $R$ exceeds a certain threshold. Moreover, we derived the corresponding scaling law for the testing error. We also demonstrated how trained abstract learners can be used for semantic compression and semantic communication.
@@
The present work is intended to inspire further thoughts on the emergence of learned skills in today’s LLMs, drawing relation to the iterative decoding procedure in communications and information theory. However, we recognize that there are still significant gaps between our theory and real-world systems. Several key factors that should be further investigated include:
@@
    Scaling laws of the number of parameters: The scaling law obtained in this paper only pertains to the size of training texts. It is unclear how the testing error scale with the number of parameters. For instance, GPT-3 is known to have more than 1,750 billion parameters, and LLMs prior to GPT-3 do not exhibit the emergence of skills. One possible explanation for the necessity of a larger number of parameters is that the capacity of a transformer, defined as the number of skills (or patterns) that can be stored in the corresponding Hopfield network [12], depends on the number of parameters in the transformer.
@@
${\cal D}$ The set of training texts
${\cal D}_{f}$ The set of domain-specific texts for training
${\cal D}_{j}$ The set of class $j$ training texts
$\tilde{\cal D}_{f}$ The set of learnable domain-specific texts
$d_{k,j}$ The average number of class $j$ texts connected to a class $k$ skill node
$d_{k}$ The mean degree of a class $k$ skill node
@@
$M(s)$ The number of edges connected to a skill node $s$
$M_{k,j}(s)$ The number of class $(k,j)$-edges connected to a class $k$ skill node
$N(t)$ The number of edges connected to a text node $t$
$N_{f}(t)$ The number of edges connected to a domain-specific text node $t$
$\tilde{N}_{f}(t)$ The number of edges connected to a learnable domain-specific text node $t$
$N_{k,j}(t)$ The number of class $(k,j)$-edges connected to a class $j$ text node
@@
${\cal S}$ The set of skills
${\cal S}_{k}$ The set of class $k$ skills
${\cal S}_{f}$ The set of domain-specific skills
$\tilde{\cal S}_{f}$ The set of learnable domain-specific skills
${\cal T}$ The set of texts composed of sequences of tokens
${\cal T}_{j}$ The set of class $j$ texts
${\cal T}_{f}$ The set of domain-specific texts
$\tilde{\cal T}_{f}$ The set of learnable domain-specific texts
@@
$\epsilon$ The probability of testing error of a randomly selected text
$\epsilon_{f}$ The probability of testing error of a randomly selected domain-specific text
@@
$\mu_{s}$ The probability that a skill node is connected to a small component via one of its edges
$\mu_{t}$ The probability that a text node is connected to a small component via one of its edges
@@
$\zeta$ The probability that a randomly selected skill is learned
$\zeta_{f}$ The probability that a randomly selected domain-specific skill is learned